OpenStudy (anonymous):

How do I solve: -8x-8y=-16 6x-9y=-108 in a matrix? I REALLY NEED HELP!!!

OpenStudy (anonymous):

first you have to find the determanate

OpenStudy (anonymous):

d= [ -8 -8] =-8(-9)-(-8)(6) = 72-(-48)= 120 [6 -9]

OpenStudy (anonymous):

I figured it out! Thank you :)

OpenStudy (anonymous):

dx=[-16 -8]= -16(-9)-(-8)(-108)= 144-864= -720 [-108 -9]

OpenStudy (anonymous):

ur welcome